the substitution method should be used to solve which system of equations
a.5x-7y=16
2x-7y=12
b.4x+3y=-5
6x-3y=2
c.x=3y+1
2x+y=7
d.2x+6y=3
3x+2y=-1

Answer by Edwin McCravy(6938)

the substitution method should be used to solve which system of 
equations 

a.5x-7y=16
2x-7y=12 
b.4x+3y=-5
6x-3y=2 
c.x=3y+1
2x+y=7 
d.2x+6y=3
3x+2y=-1


The substitution method should be used when there is a term in either of
the letters with the understood coefficient of either 1 or -1. Otherwise
the elimination method is better. 

(c) is the correct choice because it is the only one which has an x, -x, 
y or -y.  It has both an x and a y.  

Edwin